OpenStudy (lukecrayonz):

If the instructions for a problem ask you to use the smallest possible domain to completely graph two periods of y = 5 + 3 cos 2(x - π/3 ), what should be used for Xmin and Xmax?

OpenStudy (sirm3d):

set \((x- \pi/3=0)\) and \((x-\pi/3 = 4\pi)\) separately. that should give you endpoints of the two periods